OpenStudy (anonymous):

Find the leg of each isosceles right triangle when the hypotenuse is of the given measure. Given = 6sqre root two cm

OpenStudy (anonymous):

Legs are six. The sides of such a triangle have the relationship 1/1/sqrt2. You can also solve it by Pythagorean theorem.\[(6 \sqrt2)^2=2L^2 \implies 72=2L^2 \implies 6=L\]